the area of a square is 64cm2. what is the length of one side of the square?

Answer by nerdybill(7384)

the area of a square is 64cm2. what is the length of one side of the square?
Let x = length of one side of square
then
x*x = 64
x^2 = 64
x = sqrt(64)
x = 8 cm